Assume that you have a balance of $2900 on your MasterCard and that you make no more charges. Assume that MasterCard charges 12% APR and that each month you make only the minimum payment of 6% of the balance. What will the balance be after 24 months? (Round your answer to the nearest cent.)

( Mind you, It is not 846 )
Answer #1

interest in each month = opening balance * (APR / 12)

We divide by 12 as we are calculating interest for 1 month

Payment in each month = (opening balance + interest for the month) * 6%

closing balance of each month = opening balance + interest - payment

Balance after 24 months = $834.03

3 Opening closing 1 Month balance interest payment balance 2 1 $ 2,900.00 $ 29.00 $ 175.74 5 2,753.26 2 $ 2,753.26 $ 27.53 $

Opening 1 Month balance 2900 interest =B2.12%/12 =B3 12%/12 =B4.12%/12 =B5.12%/12 =B6 12%/12 =B7*12%/12 =B8*12%/12 =B9.12%/12